OpenStudy (lgbasallote):

a resistor (of unknown constant resistance) and a 0.02 farad capacitor are connected to a series with an emf of E = 100 volts. the initial charge of the capacitor is 5 coulombs. if at t=1/10 second the current flowing is 30/e amperes on the reverse direction, find the value of the resistance answer: 5 ohms

OpenStudy (hba):

this is math portion

OpenStudy (lgbasallote):

?

OpenStudy (hba):

go to the physics portion

OpenStudy (lgbasallote):

hmm you do know this is calculus right @hba

OpenStudy (hba):

soo still its a physics question

OpenStudy (lgbasallote):

nope. it's DIFFERENTIAL EQUATIONS. physics people cant help in this kind of problem. i'll be posting in the wrong group by posting there

OpenStudy (hba):

dude physics,chemistry involves mathematics that doesnt mean u post it in the math portion

OpenStudy (lgbasallote):

if you must know it follows this \[\frac{dq}{dt} + (\frac{1}{RC}) \;q = \frac{E}{R}\] ;)

OpenStudy (lgbasallote):

it's not physics

OpenStudy (lgbasallote):

it needs to be solved by calculus not physics

OpenStudy (anonymous):

go to physics section.

OpenStudy (lgbasallote):

*facepalm* @Romero

OpenStudy (anonymous):

I can't help but just so hba knows, calculus has applications in all the branches of science, chemistry, physics, biology, that'ts why calculus is so useful. To solve this physics problem, one must use calculus.

OpenStudy (lgbasallote):

but it's DIFFERENTIAL EQUATIONS

OpenStudy (anonymous):

That's not the same

OpenStudy (lgbasallote):

and i am aware of the application of calculus

OpenStudy (anonymous):

Most calculus textbooks will have examples involving the other branches..but yeah, i hope someone can help you out @igbasallote

OpenStudy (lgbasallote):

it's an application of differential equation

OpenStudy (hba):

post it in the physics portion ill give it a try

OpenStudy (anonymous):

Is this from a differential equation class or a physics class?

OpenStudy (lgbasallote):

it's a differential equations class

OpenStudy (lgbasallote):

i wouldnt post it here if it was from a physics class now would i

OpenStudy (anonymous):

Have you had physics before?

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

well not this kind

OpenStudy (lgbasallote):

i just know calculus

OpenStudy (anonymous):

Why would your professor give you a question like this. In all the real life diff eq problems that I did we had to know some previous knowledge of that field in order to set it up...

OpenStudy (lgbasallote):

i dont know..it's application of first order differential equations..that's our lesson...why wouldnt he

OpenStudy (hba):

i = dq/dt, the rate of change of charge with respect to time.

OpenStudy (lgbasallote):

fine. i'll post this in physics. but it really is just an application of firs order diff eq -_-